OpenStudy (anonymous):

Determine two pairs of polar coordinates for the point (5, -5) with 0° ≤ θ < 360°

OpenStudy (astrophysics):

Use this for polar coordinates \[x = r \cos \theta\]\[y= r \sin \theta\]\[r^2 = x^2+y^2 \implies r = \sqrt{x^2+y^2}\]

OpenStudy (anonymous):

Thank You!

OpenStudy (astrophysics):

Yw :)

OpenStudy (astrophysics):

 so you can set it up as such :-)